### v1

# fmeffects: An R Package for Forward Marginal Effects

Abstract
Forward marginal effects (FMEs) have recently been introduced as a versatile and effective model-agnostic interpretation method. They provide comprehensible and actionable model explanations in the form of: If we change $x$ by an amount $h$, what is the change in predicted outcome $\widehat{y}$? We present the R package fmeffects, the first software implementation of FMEs. The relevant theoretical background, package functionality and handling, as well as the software design and options for future extensions are discussed in this paper.
Introduction
Forward marginal effects (FMEs) (Scholbeck et al., 2022) provide simple yet accurate local model-agnostic explanations in terms of forward differences in prediction. They address questions of the form: If we change x by an amount $h$, what is the change in predicted outcome $\widehat{y}$? For instance, given a medical study where a model is trained to predict a patient’s disease risk, FMEs can tell us each patient’s individual change in predicted risk due to losing 5kg in body weight. FMEs thus provide actionable and comprehensible advice for stakeholders, including ones without expertise in machine learning. If the change in predicted risk is substantial enough, doctors may recommend a tailored exercise and nutrition regimen. To avoid interpreting the change in prediction (for instance, the change in predicted disease risk) as a linear effect, Scholbeck et al. (2022) developed a non-linearity measure (NLM) to supplement the FME. Furthermore, in applied statistics, FMEs are typically averaged to an average marginal effect (AME) for an entire data set, which serves as a global (scalar-valued) feature effect measure (Bartus, 2005). However, a single global average does not accurately represent the nuances of a non-linear model or one with complex interactions. As a remedy, Scholbeck et al. (2022) developed the conditional average marginal effect (cAME) estimator for subgroups of observations. They suggest to find subgroups with similar FMEs, for instance, by recursive partitioning (RP). FMEs therefore represent a means to explain models on a local, regional, and global level.
Contributions: We present the R package fmeffects, the first software implementation of FMEs, including the NLM and the cAME. The user interface only requires a pre-trained model and an evaluation data set. The package is designed according to modular principles and thus, various parts of the package are easily maintainable and extensible with additional functionality. This paper introduces the relevant theoretical background, explains the handling of the package with an applied interpretation use case, and discusses the software design and options for future maintenance and extensions.
Related work: Marginal effects (MEs) have a long history in applied statistics and the Stata programming language (StataCorp, 2023). Initially implemented by Bartus (2005), the margins() command is now fully integrated into Stata and provides comprehensive capabilities for various computations and visualizations (Williams, 2012). MEs are typically defined in terms of derivatives of the model w.r.t. a feature. For instance, this variant is the default approach to interpret models in econometrics (Greene, 2012). A less commonly used definition uses forward differences in prediction with varying step sizes (Scholbeck et al., 2022; Mize et al., 2019). Note that - as opposed to forward differences - derivatives are not suitable to explain piecewise constant prediction functions such as tree-based models. In recent years, MEs have gained traction in the R community. The R package margins (Leeper, 2018) was the first port of Stata’s margins() command to R. Other packages related to MEs include ggeffects (Lüdecke, 2018) and marginaleffects (Arel-Bundock, 2023). Our package, fmeffects, mainly differs from such packages in two aspects:
  • Model-agnostic: Other R packages concerned with MEs directly implement support for specific models in R, which typically does not include model classes such as random forests, support vector machines, or neural networks. In contrast, we choose a model-agnostic approach. For one, we provide support for a plethora of models trained via mlr3 (Lang et al., 2019), tidymodels (Kuhn and Wickham, 2020), and caret (Kuhn, 2008); and second, we implement an adapter pattern (Gamma et al., 1994) (see section on Design and options for extensions) that allows extensions with new model classes without substantially interfering with the remaining source code.
  • Implementing new theory surrounding FMEs: The fmeffects package is the first complete software implementation of the theory surrounding novel model-agnostic FMEs as introduced by Scholbeck et al. (2022).
Methodological background
We first provide a brief overview of the relevant theory. For an in-depth introduction, we refer the reader to the theoretical paper by Scholbeck et al. (2022). FMEs can be used for model explanations on the local, regional (also referred to as semi-global), and global level. These differ with respect to the scope of the feature space where the explanation applies. The local level explains models for single data points; the regional level for a subspace of the entire feature space (which, for randomly sampled data, translates to subgroups of observations), and the global level for the entire feature space (which, for randomly sampled data, translates to the entire data set). Increasing the “topographical” scope of the explanation requires increasing amounts of aggregations of local explanations (see for instance, the illustration by Scholbeck et al. (2020) of aggregations of local explanations to global ones for various methods). This can be problematic for non-linear models where local explanations can be highly heterogeneous due to interactions and non-linear effects estimated by the model.
Notation
[…]
Forward marginal effects
The FME can be considered a basic, local unit of interpretation. Given an observation $\bm{x}$, it tells us how the prediction changes if we move $\bm{x}$ to a different location in feature space $\bm{x}^{\ast}$. This change in feature values typically only concerns a subset of features S, where $\bm{x}^{\ast}=(\bm{x}_{S}+\bm{h}_{S},\bm{x_{-S}})$:
$\text{FME}_{\bm{x},\bm{h}_{S}}$ $=\widehat{f}(\bm{x_{S}}+\bm{h}_{S},\bm{x_{-S}})-\widehat{f}(\bm{x})\quad\quad\text{for continuous features $\bm{x_{S}}$}$
For categorical features, this shift is only possible for observations whose observed values differ from the reference location. The observation-specific categorical FME with reference category $c_{j}$ corresponds to:
$\text{FME}_{\bm{x},c_{j}}=\widehat{f}(c_{j},\bm{x_{-j}})-\widehat{f}(\bm{x})\quad\quad\text{for categorical $x_{j}$}$
Non-linearity measure
For continuous features, we can consider $\bm{x}_{S}+\bm{h}_{S}$ a continuous transition of feature values. The associated change in prediction may be misinterpreted as a piecewise linear effect, which is counteracted by the non-linearity measure (NLM). The NLM corresponds to a continuous coefficient of determination $R^{2}$ between prediction function and the linear secant that intersects $\bm{x}$ and $\bm{x}^{\ast}$ (see Fig. 1). We first parameterize this continuous transition through the feature space as a fraction $t\in[0,1]$ of the multivariate step size $\bm{h}_{S}$:
$\gamma(t)=\bm{x}+t\cdot(\bm{h}_{S},0,\dots,0)$
Given the linear secant $g_{\bm{x},\bm{h}_{S}}(\gamma(t))$, the NLM corresponds to:
[…]
Fig. 2 illustrates the setting for multivariate feature changes. The NLM can be approximated via numerical integration, for instance via Simpson’s rule.
Figure 1: Illustration by Scholbeck et al. (2022) of a univariate FME (blue bar) given the prediction function (yellow) and linear secant (green). The NLM indicates how well the secant can explain the prediction function - which is inversely proportional to the integral visualized in red - in relation to how well the most uninformative baseline model - the average prediction - can explain the prediction function.
Figure 2: Illustration of the multivariate NLM by Scholbeck et al. (2022). Left: An exemplary bivariate prediction function and two points to compute an FME. Our observation is located at point (-5, -5), and we create the shortest path through the feature space to reach point (5, 5), which consists of directly proportional changes in both features. Above the path, we see the linear secant and the non-linear prediction function. Right: The multivariate change in feature values can be parameterized as a percentage $t$ of the step size $\bm{h}_{S}$. As a result, the deviation between prediction function and linear secant, as well as the deviation between the prediction function and mean prediction both correspond to a line integral which can be numerically approximated, for instance via Simpson’s rule.
The NLM indicates how well the linear secant is able to explain the prediction function, compared to the baseline model of using the mean prediction. A value of 1 indicates perfect linearity, where the linear secant is identical to the prediction function. For a value of 0, the mean prediction is able to explain the prediction function as equally well as the secant. For negative values, the mean prediction better explains the prediction function than the linear secant (severe non-linearity).
It is, therefore, easiest to interpret FMEs with NLM values close to 1. Although every FME always represents the exact change in prediction, an FME with a low NLM value does not fully describe the behavior of the model in that specific locality. In contrast, an FME with an NLM close to 1 is a sufficient descriptor of the (linear) model behavior. In other words, the NLM serves as an auxiliary diagnostic tool, indicating trust in how well the FME describes the local change in prediction.
Conditional average marginal effect
In order to receive a global model explanation akin to a beta coefficient in linear models, local FMEs can be averaged to a global average marginal effect (AME). However, this does not accurately represent the nuances of a non-linear prediction function. Given a data set $\mathcal{D}$, the conditional average marginal effect (cAME) estimator applies to a subgroup of observations $\mathcal{D}_{[\;]}$:
$\;\text{cAME}_{\mathcal{D}_{[\;]},\bm{h}_{S}}$ $=\widehat{\mathbb{E}_{\bm{X}_{[\;]}}\left[\text{FME}_{\bm{X}_{[\;]},\bm{h}_{S}}\right]}$
$=\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{f}\left(\bm{x_{S}}^{(i)}+\bm{h}_{S},\bm{x_{-S}}^{(i)}\right)-\widehat{f}\left(\bm{x^{(i)}}\right)\right)\,\forall i:\bm{x^{(i)}}\in\mathcal{D}_{[\;]}$
Although this estimator can be applied to arbitrary subgroups, an interpretation task demands searching for subgroups with homogeneous local model explanations. Desiderata for such subgroups include within-group effect homogeneity, between-group effect heterogeneity, full segmentation, non-congruence, confidence, and stability. A viable option to partition $\mathcal{D}$ is to run a recursive partitioning algorithm on $\mathcal{D}$ with FMEs as the target which aims at finding subgroups where FMEs are similar. For instance, in fmeffects, both rpart (Therneau and Atkinson, 2019) and ctree() from partykit (Hothorn and Zeileis, 2015) are supported to find subgroups.
User interface and package handling
Local explanations
The fme() function is the central user interface, which - apart from certain parameters - only requires a pre-trained model and a data set (see section Design and options for extensions for details). Further arguments specify the target, feature(s), step size(s), whether to compute NLM values for each FME, and specification of an extrapolation detection method [^3] Model extrapolations are predictions in areas with a low density of training observations. Scholbeck et al. (2022) provide a short introduction to this topic.. The fme() function initiates construction of and computations for a ForwardMarginalEffects object without requiring the user to know R6 (Chang, 2021) syntax.
For this use case, we train a random forest from the ranger package (Wright and Ziegler, 2017) on the bike sharing data set (Fanaee-T, 2013) using mlr3. Note that models trained via tidymodels and caret are also supported. Our goal is to predict and explain the daily bike rental demand in Washington D.C. based on features such as the outside temperature, wind speed, or humidity. We first train the model:
> library(fmeffects)> data(bikes, package = "fmeffects")> library(mlr3verse)> forest = lrn("regr.ranger")> task = as_task_regr(x = bikes, id = "bikes", target = "count")> forest\$train(task)
Then, we simply pass the trained model, evaluation data, and remaining parameters to the fme() function. It returns a ForwardMarginalEffects object which can be analyzed via summary() and visualized via plot() (see Fig. 3). Here, the outside temperature is raised by 5 degrees Celsius ceteris paribus. Each dot represents a single instance. The arrow above the horizontal axis indicates the step size and direction of the feature change in relation to the feature range. The horizontal value of each instance indicates its observed feature value in the provided data set; the vertical axis value is the corresponding FME.
Let us single out the observation with the largest associated FME (brown colored and enlarged in Fig. 3). This observation corresponds to a single day with a recorded temperature of 13.12 degrees Celsius. Increasing the temperature by 5 degrees Celsius on this particular day results in an additional 125 predicted bike rentals. We plot such model explanations for the entire data set and average FMEs to receive a global model explanation. The AME - the global average of FMEs - is 14.5: An increase in temperature by 5 degrees Celsius results in an average change in predicted daily bike rentals by 14.5.
> effects = fme(+   model = forest,+   data = bikes,+   target = "count",+   feature = "temp",+   step.size = 5,+   ep.method = "envelope")> summary(effects)
Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5Extrapolation point detection:  envelope, EPs: 81 of 727 obs. (11 %)Average Marginal Effect (AME):  14.4844
> plot(effects)
Figure 3: Plot of univariate FMEs for feature ‘temp’ and step size 5. Each dot represents an observation. The horizontal value represents the observed value of feature ‘temp’. We move each observation’s ‘temp’ value according to the arrow’s direction and length (in this case, 5 units to the right). The respective FME value is indicated by the vertical value of each point. The horizontal bar indicates the AME (the global FME average). An additional smoothing function facilitates interpretations by modeling an approximate pattern of FMEs across the feature range. We single out one observation (brown colored and enlarged) to demonstrate a local interpretation.
Let us take a moment to compare the FME plot with the individual conditional expectation (ICE) (Goldstein et al., 2015) and partial dependence (PD) (Friedman, 2001) plot generated by the R package iml (Molnar et al., 2018). The ICE and PD are one of the most popular and established model-agnostic ways to interpret predictive models (Molnar, 2022). The ICE is a local model explanation and represents the prediction for an observation where only the features of interest are varied (in this case only ‘temp’). The PD is the average of ICEs (in the univariate case, the vertical average) and indicates the global, average prediction when a subset of features is varied for all observations. From Fig. 4, we gather that in contrast to FMEs, the ICE and PD plot is more difficult to interpret. Although we can see a rough trajectory of the feature influence on local and average predictions, it is difficult to pinpoint the exact effects of changing ‘temp’ on the prediction for single observations. Furthermore, ICE curves are more likely to be subject to model extrapolations, a result from predicting in areas where the model was not trained on a sufficient amount of data.
Figure 4: An ICE and PD plot for feature ‘temp’ generated by the R package iml. Each black curve (an ICE) represents predictions for a single instance while only feature ‘temp’ varies. The yellow curve (the PD) is the vertical average of ICEs and represents the average, isolated influence of ‘temp’ on predictions. As opposed to FMEs, this plot is substantially more difficult to interpret.
In many applications, we are interested in interactions of features on the prediction. We diagnosed a univariate effect of ‘temp’ on the predicted amount of bike rentals, but effects may be enhanced or mitigated when other features are varied at the same time. For instance, humidity has a considerable impact on the effects of heat (Russo et al., 2017). We evaluate an increase in temperature by 5 degrees Celsius and a simultaneous drop in humidity by 10 percentage points (see Fig. 5). For a bivariate change in feature values, the two arrows depict the direction and magnitude of the feature change in the respective variable. Each point is located at the observation’s observed feature values in the provided data set. Its color indicates the FME associated with the bivariate feature change. An increase in the outside temperature by 5 degrees Celsius and a drop in humidity by 10 percentage points is associated with an AME of 19.14. The isolated univariate effect of increasing the outside temperature is boosted by the simultaneous drop in humidity.
> effects.bivariate = fme(model = forest,+               data = bikes,+              target = "count",+              feature = c("temp", "humidity"),+              step.size = c(5, -0.1),+              ep.method = "envelope")> summary(effects.bivariate)
Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5  humidity, -0.1Extrapolation point detection:  envelope, EPs: 82 of 727 obs. (11 %)Average Marginal Effect (AME):  19.1418
> plot(effects.bivariate)
Figure 5: Visualizing bivariate FMEs for an increase in ‘temp’ by 5 degrees Celsius and a simultaneous drop in ‘humidity’ by 10 percentage points. FMEs are highly heterogeneous. We can see mostly positive effects, especially for observations with combinations of lower ‘temp’ and ‘humidity’ values. For various really warm days, the model even predicts a negative effect.
Let us repeat the same procedure as for univariate feature changes and compare the FME plot to an alternative option, the bivariate PD plot (see Fig. 6). As opposed to the novel visualization with FMEs, the PD plot only visualizes the average, global effect of changing both features on the predicted amount of bike rentals, and it does not inform about the distribution of observed feature values, thus not allowing us to evaluate effects of increasing one feature and decreasing another at the same time. Furthermore, the PD plot - like its univariate counterpart - is more likely to be subject to model extrapolations due to predicting for many feature value combinations the model has not seen during training.
Figure 6: A bivariate PD plot (created via the R package iml), one of the few available visualization tools for interactions. We can see the global interaction between ‘temp’ and ‘humidity’ on the predicted amount of bike rentals. Plugging in larger values for both features ceteris paribus results in more predicted bike rentals on average. As opposed to bivariate FMEs, we cannot investigate multiple local effects, nor can we see the actual distribution of observed feature values, and as a result, we cannot evaluate the effects of increasing one feature and decreasing another at the same time.
Let us now proceed to investigations of non-linearity. Although some sense of non-linearity can be derived from ICE curves (see Fig. 4), it is hard to quantify, and summarizing non-linearity for an ICE with a large variation in the feature of interest would be somewhat meaningless. Furthermore, for bivariate or higher-dimensional changes in feature values, we lose any option for visual diagnoses of non-linearity. In contrast, the NLM can be computed for FMEs with continuous step sizes, regardless of their dimensionality:
> effects.bivariate.nlm = fme(model = forest,+                       data = bikes,+                       target = "count",+                       feature = c("temp", "humidity"),+                       step.size = c(5, -0.1),+                       ep.method = "envelope",+                       compute.nlm = TRUE)> plot(effects.bivariate.nlm, with.nlm = TRUE)
Figure 7: Adding NLM computations to the FME plot. Each dot in the left and right plot represents a single observation and its respective FME and NLM value. We can see both linear (bright, yellow colored NLM) and non-linear FMEs (dark, blue colored NLM).
Fig. 7 simply contrasts each observation’s NLM value with its FME value. In this case, we can see both non-linear FMEs (darker, blue colored NLM values) and linear FMEs (brighter, yellow colored NLM values). This is not surprising, considering that the bike sharing data consists of many features, each being capable to influence the effect of changes in ‘temp’ and ‘humidity’ on the prediction.
So far, we have only evaluated changes in continuous features. In many applications, we are often concerned with switching categories of categorical features, a way of counterfactual thinking inherent to the human thought process. Note that in spite of the allure of switching categories of categorical features, one needs to be aware of potential model extrapolations. To illustrate this, we switch each non-rainy day’s precipitation status to rainfall. Rainfall has an average isolated effect of lowering daily rentals by 57 bikes (see Fig. 8).
> effects.categ = fme(model = forest,+                data = bikes,+                target = "count",+                feature = "weather",+                step.size = "rain")> summary(effects.categ)Forward Marginal Effects ObjectStep type:  categoricalFeature & reference category:  weather, rainExtrapolation point detection:  none, EPs: 0 of 657 obs. (0 %)Average Marginal Effect (AME):  -57.3322> plot(effects.categ)
Figure 8: Distribution of categorical FMEs resulting from switching each non-rainy day’s precipitation status to rain. On average, rainfall lowers predicted bike rentals by 57 bikes per day.
Regional explanations
In our examples, we can see highly heterogeneous local effects. The more heterogeneous FMEs are, the less information the AME carries. In many practical applications, we are interested in compactly describing the behavior of the predictive model across the feature space, akin to a beta coefficient in a linear model. This is where regional explanations come into play. We aim to find subgroups with more homogeneous FME values, thereby describing the behavior of the model not in terms of a global average, but in terms of multiple regional averages (cAMEs).
In fmeffects, this can be achieved by further processing the ForwardMarginalEffects object containing FMEs (and optionally NLM values) using the came() function. This returns a Partitioning object (in this case, an object of the class "PartitioningCTREE", a subclass of the abstract class "Partitioning", see later section on Design and options for extensions).
For the univariate change in temperature by 5 degrees Celsius, we decide to search for exactly 3 subgroups [^4] This value is to be set by the user depending on how many regional explanations are to be found. How many subgroups can be found depends on the data and predictive model. (for a description of this algorithm, see the following section on Design and options for extensions). A summary of the created object informs us about the number of observations, cAME, and standard deviation (SD) of FMEs inside the root node and leaf nodes (the found subgroups). We succeed in finding subgroups with lower SDs while maintaining an appropriate sample size. The root node SD of 22.5 can be successfully split down to 21.5, 8.5, and 17.7 inside the subgroups. By visualizing the tree, we can see how the data was partitioned. Lower temperatures are associated with larger increases in predicted bike rentals when the outside temperature is raised. For outside temperatures lower than 17 degrees Celsius and working days, we can observe the largest cAME of 32 additional bike rentals per day and a cAME of 11 additional daily bike rentals for weekends. For outside temperatures above 17 degrees Celsius, there is only a slight effect of 1.7 additional bikes per day:
> subspaces = came(effects = effects, number.partitions = 3)> summary(subspaces)
PartitioningCtree of an FME objectMethod:  partitions = 3   n      cAME   SD(FME) 646 14.484353 22.499156 * 230 32.265791 21.546565 130 11.172507  8.554322 286  1.689979 17.749573---* root node (non-partitioned)AME (Global): 14.4844
> plot(subspaces)
Figure 9: Finding subgroups of observations with a decision tree. Each leaf node visualizes one subgroup, the number of observations, the cAME, and the SD of FMEs indicating FME heterogeneity.
Global explanations
When to search for regional explanations thus depends on the heterogeneity of local effects, which can be summarized for the entire model via the ame() function. It uses a default step size of 1 or 0.01 for small feature ranges. For categorical FMEs, it uses every observed category as a reference category. The summary() function prints a compact model summary of each feature, a default step size, the AME, the SD of FMEs, 25% and 75% quantiles of FMEs, as well as the number of observations left after excluding extrapolation points. A large dispersion indicates heterogeneity of FMEs and thus, a small fidelity of the AME and possible benefits from searching for subgroups with varying cAMEs. A different workflow can, therefore, also consist of starting with the table generated by ame() and deciding which feature effects can be described by AMEs, which might be better describable by subgroups and cAMEs, and - if this has been without success - resorting to local model explanations. Recall our example from the previous section on Regional explanations where we split FMEs associated with increasing temperature by 5 degrees Celsius. From the ame() summary, we gather that ‘temp’ has a relatively large SD in relation to its AME, and the interquartile range indicates a large spread of FMEs from -0.47 in the 25% quantile up to 4.74 in the 75% quantile, thus making it a promising candidate to find subgroups with more homogeneous FMEs.
> ame.results = ame(model = forest, data = bikes, target = "count")> summary(ame.results)
Model Summary Using Average Marginal Effects:      feature step size       AME      SD      0.25      0.75   n1      season    spring  -29.3415 30.6189  -38.2318   -6.3963 5482      season    summer    0.5742 21.4782   -8.3499   10.5607 5433      season      fall      9.97 27.4275   -1.9819    30.745 5394      season    winter   13.9861 23.6806    1.0266   23.6969 5515        year         0  -98.6521 67.2816 -156.4247    -19.63 3646        year         1   96.6155 60.1458     22.91  148.1915 3637       month         1    3.9954 13.4096   -1.2981    7.2207 7278     holiday     False   -1.5799 20.5126  -12.5858   14.0711  219     holiday      True  -12.0674 24.2779  -29.8895    6.4375 70610    weekday       Sat  -54.4458 49.1836  -85.8245  -15.5722 62211    weekday       Sun  -78.6803  51.462 -113.7695  -31.2021 62212    weekday       Mon   10.7612 28.3258   -8.7131   30.5779 62313    weekday       Tue   17.9565 24.7443    0.9637   33.5499 62514    weekday       Wed    20.721 22.6949    1.5863   34.3365 62315    weekday       Thu   19.7423 23.9554   -0.2864   34.7076 62416    weekday       Fri    1.6724 35.8582  -23.9194   29.7121 62317 workingday     False -204.2023 90.1427 -256.0286 -141.4579 49618 workingday      True  165.2412 64.8985  123.8688   213.351 23119    weather     clear   25.9084  42.223    3.6162   22.3112 28420    weather     misty    3.3301 33.2101   -9.0864    0.8842 51321    weather      rain  -57.3322 55.5617  -95.7276   -3.8571 65722       temp         1    2.3959  7.2201   -0.4717    4.7369 72723   humidity      0.01   -0.1681  2.4691   -0.2999    0.4329 72724  windspeed         1    0.0198  2.2861   -0.2156    0.2126 727
Design and options for extensions
The fmeffects package is built on a modular design for improved maintainability and future extensions. Fig. 10 provides a visual overview of the core design. Major emphasis is given to the strategy and adapter design patterns (Gamma et al., 1994). Simply put, the strategy pattern decouples source code for algorithm selection at runtime into separate classes. We repeatedly implement this pattern throughout the package by creating abstract classes whose subclasses implement specific functionalities. The adapter design pattern (also referred to as a “wrapper”) creates an interface for communication between two classes.
Figure 10: Design overview of the fmeffects package, including methods that implement the main functionality of each class. Classes may contain more methods than depicted. Blue boxes indicate wrapper functions to instantiate objects of the respective class.
  • "Predictor": An abstract class that implements the adapter pattern to accommodate future implementations of storing a predictive model. "PredictorMLR3", "PredictorParsnip", and "PredictorCaret" are subclasses that store an mlr3, parsnip (Kuhn and Vaughan, 2023) (part of tidymodels), or caret model object. This allows users of fmeffects to use numerous predictive models such as random forests, gradient boosting, support vector machines, or neural networks. The package can be extended with novel model types by implementing a new subclass that stores the model, data, target, and is able to return predictions.
  • "AverageMarginalEffects": A class to compute AMEs for each feature in the data (or a subset thereof). Internally, a new "ForwardMarginalEffects" object is used to compute FMEs and aggregate them. For convenience, we implement a wrapper function ame() to facilitate object creation and to initiate computations without requiring user input in the form of R6 syntax.
  • "ForwardMarginalEffects": The centerpiece class of the package. It keeps access to a Predictor, stores important information to create FMEs, and after computations are completed, stores results and gives access to visualization methods. For convenience, the wrapper function fme() can be used.
  • "FMEPlot": An abstract class for code decoupling of different plot categories into distinct classes. Subclasses include "FMEPlotUnivariate", "FMEPlotBivariate", "FMEPlotCategorical".
  • "ExtrapolationDetector": Identifies (and excludes) extrapolation points. The current implementation supports the method “envelope”, which excludes points outside of the multivariate envelope of the training data (Scholbeck et al., 2022).
  • "NonLinearityMeasure": For the NLM, we need to approximate three line integrals, for instance via Simpson’s 3/8 rule. The general definition for a univariate function f(x) and integration interval $[a,b]$ corresponds to:
    $\int_{a}^{b}f(x)\approx\frac{b-a}{8}\left[f(a)+3f\left(\frac{2a+b}{3}\right)+3f\left(\frac{a+2b}{3}\right)+f(b)\right]$ (1)
    We make use of a composite Simpson rule, which divides up the interval $\left[a,b\right]$ into $n$ subintervals of equal size and approximates each subinterval with Eq. (1).
  • "Partitioning": An abstract class, allowing various implementations of finding subgroups for cAMEs. For convenience, the wrapper function came() can be used. The current implementation supports RP via the rpart and partykit (CTREE algorithm) packages (classes "PartitioningRPart" and "PartitioningCTREE"). We believe there are two criteria that should guide this process:
    • FME homogeneity inside each subgroup.
    • The number of subgroups.
    A low number of subgroups is generally preferred. In certain applications, we may want to search for a predefined number of subgroups, akin to the search for a predefined number of clusters in clustering problems. Many RP algorithms do not support the search for a number of subgroups, which is where the "Pruner" class comes into play.
  • "Pruner": To receive a predefined number of subgroups for arbitrary RP algorithms, we follow a two-stage process: Grow a large tree by tweaking tree-specific hyperparameters and then prune it back to receive the desired number of subgroups. A "Partitioning" subclass is implemented such that it can first grow a large tree, for instance, with a low complexity parameter for rpart. Then, "Pruner" iteratively computes the SD of FMEs for each parent node of the current terminal nodes and removes all terminal nodes of the parent with the lowest SD.
  • "PartitioningPlot": Decouples visualizations of separating $\mathcal{D}$ into subgroups from the concrete implementation of the "Partitioning" subclass. Here, we make use of a dependency on partykit for a tree data structure. This allows visualizations of any partitioning with the same methods. The package ggparty (Borkovec and Madin, 2019) creates tree figures that illustrate the partitioning, descriptive statistics for each terminal node, and histograms of FMEs (and optionally NLM values).
Conclusion
This paper introduces the R package fmeffects, the first software implementation of FMEs. We showcase the package functionality with an applied use case and discuss design choices and implications for future extensions. FMEs are a novel and versatile method for model-agnostic interpretations. They give us explanations in the form of: If we change $x$ by an amount $h$, what is the change in predicted outcome $\widehat{y}$? FMEs provide comprehensible and actionable advice for stakeholders, including ones without machine learning expertise. We therefore hope that this package will work towards a more widespread adaption and usage of FMEs in practice.

### v2

@@
 Abstract
-Forward marginal effects (FMEs) have recently been introduced as a versatile and effective model-agnostic interpretation method.
-They provide comprehensible and actionable model explanations in the form of: If we change $x$ by an amount $h$, what is the change in predicted outcome $\widehat{y}$?
-We present the R package fmeffects, the first software implementation of FMEs.
+Forward marginal effects have recently been introduced as a versatile and effective model-agnostic interpretation method particularly suited for non-linear and non-parametric prediction models.
+They provide comprehensible model explanations of the form: if we change feature values by a pre-specified step size, what is the change in the predicted outcome?
+We present the R package fmeffects, the first software implementation of the theory surrounding forward marginal effects.
 The relevant theoretical background, package functionality and handling, as well as the software design and options for future extensions are discussed in this paper.
 Introduction
-Forward marginal effects (FMEs) (Scholbeck et al., 2022) provide simple yet accurate local model-agnostic explanations in terms of forward differences in prediction.
-They address questions of the form: If we change x by an amount $h$, what is the change in predicted outcome $\widehat{y}$?
-For instance, given a medical study where a model is trained to predict a patient’s disease risk, FMEs can tell us each patient’s individual change in predicted risk due to losing 5kg in body weight.
-FMEs thus provide actionable and comprehensible advice for stakeholders, including ones without expertise in machine learning.
-If the change in predicted risk is substantial enough, doctors may recommend a tailored exercise and nutrition regimen.
-To avoid interpreting the change in prediction (for instance, the change in predicted disease risk) as a linear effect, Scholbeck et al.
-(2022) developed a non-linearity measure (NLM) to supplement the FME.
-Furthermore, in applied statistics, FMEs are typically averaged to an average marginal effect (AME) for an entire data set, which serves as a global (scalar-valued) feature effect measure (Bartus, 2005).
-However, a single global average does not accurately represent the nuances of a non-linear model or one with complex interactions.
-As a remedy, Scholbeck et al.
-(2022) developed the conditional average marginal effect (cAME) estimator for subgroups of observations.
-They suggest to find subgroups with similar FMEs, for instance, by recursive partitioning (RP).
-FMEs therefore represent a means to explain models on a local, regional, and global level.
-Contributions: We present the R package fmeffects, the first software implementation of FMEs, including the NLM and the cAME.
+A growing number of disciplines are adopting black box machine learning (ML) models to make predictions, including medicine (Rajkomar et al., 2019; Boulesteix et al., 2020), psychology (Dwyer et al., 2018), economics (Mullainathan and Spiess, 2017; Athey and Imbens, 2019), or the earth sciences (Dueben and Bauer, 2018).
+Although one can often observe a superior predictive performance of black box models (such as neural networks, gradient boosting, random forests, or support vector machines) over intrinsically interpretable models (such as generalized linear or additive models), their lack of transparency or interpretability is considered a major drawback (Breiman, 2001).
+This has been a major driver in the development of model-agnostic explanation techniques, which are often referred to by the umbrella terms of interpretable ML (Molnar, 2022) or explainable artificial intelligence (Kamath and Liu, 2021).
+Marginal effects (MEs) (Williams, 2012) have been a mainstay of model interpretations in many applied fields such as econometrics (Greene, 2019), psychology (McCabe et al., 2022), or medical research (Onukwugha et al., 2015).
+MEs explain the effect of features on the predicted outcome in terms of derivatives w.r.t.
+a feature or forward differences in prediction.
+They are typically averaged to an average marginal effect (AME) for an entire data set, which serves as a global (scalar-valued) feature effect measure (Bartus, 2005).
+To explain feature effects for non-linear models, Scholbeck et al.
+(2024) introduced a unified definition of forward marginal effects (FMEs), a non-linearity measure (NLM) for FMEs, and the conditional average marginal effect (cAME).
+The NLM is an auxiliary model diagnostic to avoid interpreting local changes in prediction as linear effects.
+The cAME aims to describe the model via regional FME averages for subgroups with similar FMEs, which can, for instance, be found by recursive partitioning (RP).
+FMEs, therefore, represent a means to explain models on a local, regional, and global level.
+Contributions: We present the R package fmeffects, the first software implementation of the theory surrounding FMEs, including the NLM and the cAME.
 The user interface only requires a pre-trained model and an evaluation data set.
-The package is designed according to modular principles and thus, various parts of the package are easily maintainable and extensible with additional functionality.
-This paper introduces the relevant theoretical background, explains the handling of the package with an applied interpretation use case, and discusses the software design and options for future maintenance and extensions.
-Related work: Marginal effects (MEs) have a long history in applied statistics and the Stata programming language (StataCorp, 2023).
-Initially implemented by Bartus (2005), the margins() command is now fully integrated into Stata and provides comprehensive capabilities for various computations and visualizations (Williams, 2012).
-MEs are typically defined in terms of derivatives of the model w.r.t.
-a feature.
-For instance, this variant is the default approach to interpret models in econometrics (Greene, 2012).
-A less commonly used definition uses forward differences in prediction with varying step sizes (Scholbeck et al., 2022; Mize et al., 2019).
-Note that - as opposed to forward differences - derivatives are not suitable to explain piecewise constant prediction functions such as tree-based models.
-In recent years, MEs have gained traction in the R community.
-The R package margins (Leeper, 2018) was the first port of Stata’s margins() command to R.
-Other packages related to MEs include ggeffects (Lüdecke, 2018) and marginaleffects (Arel-Bundock, 2023).
-Our package, fmeffects, mainly differs from such packages in two aspects:
-  • Model-agnostic: Other R packages concerned with MEs directly implement support for specific models in R, which typically does not include model classes such as random forests, support vector machines, or neural networks.
-In contrast, we choose a model-agnostic approach.
-For one, we provide support for a plethora of models trained via mlr3 (Lang et al., 2019), tidymodels (Kuhn and Wickham, 2020), and caret (Kuhn, 2008); and second, we implement an adapter pattern (Gamma et al., 1994) (see section on Design and options for extensions) that allows extensions with new model classes without substantially interfering with the remaining source code.
-  • Implementing new theory surrounding FMEs: The fmeffects package is the first complete software implementation of the theory surrounding novel model-agnostic FMEs as introduced by Scholbeck et al.
-(2022).
-Methodological background
-We first provide a brief overview of the relevant theory.
-For an in-depth introduction, we refer the reader to the theoretical paper by Scholbeck et al.
-(2022).
+The package is designed according to modular principles, making it simple to maintain and extend.
+This paper introduces the relevant theoretical background of FMEs, demonstrates the usage of the package in the context of a practical use case, and explains the software design.
+Background on forward marginal effects
 FMEs can be used for model explanations on the local, regional (also referred to as semi-global), and global level.
-These differ with respect to the scope of the feature space where the explanation applies.
-The local level explains models for single data points; the regional level for a subspace of the entire feature space (which, for randomly sampled data, translates to subgroups of observations), and the global level for the entire feature space (which, for randomly sampled data, translates to the entire data set).
-Increasing the “topographical” scope of the explanation requires increasing amounts of aggregations of local explanations (see for instance, the illustration by Scholbeck et al.
+These differ with respect to the region of the feature space that the explanation refers to.
+The local level explains a model/prediction for single observations, the regional level for a certain subspace (or subgroups of observations), and the global level for the entire feature space.
+Increasing the scope of the explanation requires increasing amounts of aggregations of local explanations (see the illustration by Scholbeck et al.
 (2020) of aggregations of local explanations to global ones for various methods).
-This can be problematic for non-linear models where local explanations can be highly heterogeneous due to interactions and non-linear effects estimated by the model.
+This can be problematic for non-parametric models where local explanations can be highly heterogeneous due to non-linear effects or interactions.
 Notation
-[…]
+Let $\widehat{f}:\mathcal{X}\rightarrow\mathbb{R}$ be the prediction function of a learned model where $\mathcal{X}\subset\mathbb{R}^{p}$ denotes the feature space.
+While our definition naturally covers regression models, for classification models, we assume that $\widehat{f}$ returns the score or probability for a predefined class of interest.
+A subspace of the feature space is denoted by $\mathcal{X}_{[\;]}\subseteq\mathcal{X}$.
+The random feature vector is denoted by [^1] Bold letters denote vectors.
+$\bm{X}=(X_{1},\dots,X_{p})$.
+Observations are denoted by $\bm{x}=(x_{1},\dots,x_{p})\in\mathcal{X}$.
+A set of feature indices is denoted by $S\subseteq\{1,\dots,p\}$.
+We often index (random) vectors as $\bm{x}_{S}$ or $\bm{X}_{S}$.
+We denote set complements by $-S=\{1,\,\dots\,,\,p\}\;\setminus\;S$.
+With slight abuse of notation, we represent the partitioning of a vector into two arbitrary but disjoint groups by $\bm{x}=(\bm{x}_{S},\bm{x}_{-S})$, regardless of the order of features.
+For a single feature of interest, the set $S$ is replaced by an integer index $j$.
+We usually assume an evaluation data set $\mathcal{D}=\left(\bm{x}^{(i)}\right)_{i=1}^{n}$, with $\bm{x}^{(i)}\in\mathcal{X}$, which may consist of both training and test data.
 Forward marginal effects
 The FME can be considered a basic, local unit of interpretation.
-Given an observation $\bm{x}$, it tells us how the prediction changes if we move $\bm{x}$ to a different location in feature space $\bm{x}^{\ast}$.
-This change in feature values typically only concerns a subset of features S, where $\bm{x}^{\ast}=(\bm{x}_{S}+\bm{h}_{S},\bm{x_{-S}})$:
+Given an observation $\bm{x}$, it tells us how the prediction changes if we change a subset of feature values $\bm{x}_{S}$ by a vector of step sizes $\bm{h}_{S}$.
 $\text{FME}_{\bm{x},\bm{h}_{S}}$ $=\widehat{f}(\bm{x}_{S}+\bm{h}_{S},\bm{x}_{-S})-\widehat{f}(\bm{x})\quad\quad\text{for continuous features $\bm{x}_{S}$}$
-For categorical features, this shift is only possible for observations whose observed values differ from the reference location.
-The observation-specific categorical FME with reference category $c_{j}$ corresponds to:
+Scholbeck et al.
+(2024) introduced an observation-specific categorical FME whose definition is congruent with the FME for continuous features.
+The categorical FME corresponds to the change in prediction when replacing $x_{j}$ by the reference category $c_{j}$:
 $\text{FME}_{\bm{x},c_{j}}=\widehat{f}(c_{j},\bm{x}_{-j})-\widehat{f}(\bm{x})\quad\quad\text{for categorical $x_{j}$}$
+Note that this definition of a categorical ME differs from the one that is typically found in fields like econometrics (Williams, 2012), where we set $x_{j}$ to a reference category for all observations and then record the change in prediction resulting from changing the reference category to another category.
+Furthermore, it is common to globally average MEs to an average marginal effect (AME) to estimate the expected local effect. For FMEs, this corresponds to:
+$\;\text{AME}_{\mathcal{D},\bm{h}_{S}}$ $=\widehat{\mathbb{E}_{\bm{X}}\left[\text{FME}_{\bm{X},\bm{h}_{S}}\right]}$
+$=\frac{1}{n}\sum_{i=1}^{n}\left[\widehat{f}\left(\bm{x}_{S}^{(i)}+\bm{h}_{S},\bm{x}_{-S}^{(i)}\right)-\widehat{f}\left(\bm{x}^{(i)}\right)\right]$ (1)
+Note that for categorical feature changes and observations where $x_{j}=c_{j}$, the FME equals 0.
+In the fmeffects package, the categorical AME only consists of observations whose observed feature value differs from the selected category.
+This approach is motivated by our goal to explain the effects of changing feature values on the predicted outcome.
+For instance, in Fig.
+11, we demonstrate the effect of rainfall on the daily number of bike rentals in Washington D.C.
+by switching each non-rainy day’s precipitation status to rainfall.
+Considering all observations, including rainy days, would obfuscate the interpretation we desire from our model.
+However, it is important to remember that every AME comprises a different set of points.
+Step size selection
+The selection of step sizes is determined by contextual and data-related considerations (Scholbeck et al., 2024).
+First, the FME allows us to investigate the model according to specific research questions.
+For instance, we might be interested in the effects of a specific change in a patient’s body weight on the predicted individual disease risk.
+Often, we are interested in an interpretable or intuitive step size.
+In the case of body weight, typically expressed in kilograms, we could use a 1kg change (for instance, instead of 1g) as a natural increment.
+Without contextual information, we could use a unit change as a reasonable default; or dispersion-based measures such as one standard deviation, percentages of the interquartile range, or the mean/median absolute deviation.
 Non-linearity measure
 For continuous features, we can consider $\bm{x}_{S}+\bm{h}_{S}$ a continuous transition of feature values.
-The associated change in prediction may be misinterpreted as a piecewise linear effect, which is counteracted by the non-linearity measure (NLM).
-The NLM corresponds to a continuous coefficient of determination $R^{2}$ between prediction function and the linear secant that intersects $\bm{x}$ and $\bm{x}^{\ast}$ (see Fig.
+The associated change in prediction may be misinterpreted as a linear effect.
+This is counteracted by the NLM, which corresponds to a continuous coefficient of determination $R^{2}$ between the prediction function and the linear secant that intersects $\bm{x}$ and $(\bm{x}_{S}+\bm{h}_{S},\bm{x}_{-S})$ (see Fig.
 1).
-We first parameterize this continuous transition through the feature space as a fraction $t\in[0,1]$ of the multivariate step size $\bm{h}_{S}$:
-$\gamma(t)=\bm{x}+t\cdot(\bm{h}_{S},0,\dots,0)$
-Given the linear secant $g_{\bm{x},\bm{h}_{S}}(\gamma(t))$, the NLM corresponds to:
+The continuous transition through the feature space is first parameterized as a fraction $t\in[0,1]$ of the multivariate step size $\bm{h}_{S}$:
+$\gamma(t)=\begin{pmatrix}x_{1}\\ \vdots\\ x_{p}\end{pmatrix}+t\cdot\begin{pmatrix}h_{1}\\ \vdots\\ h_{s}\\ 0\\ \vdots\\ 0\end{pmatrix}\quad,\quad t\in[0,1]$
+The value of the linear secant $g_{\bm{x},\bm{h}_{S}}(t)$ corresponds to:
 […]
-Fig. 2 illustrates the setting for multivariate feature changes. The NLM can be approximated via numerical integration, for instance via Simpson’s rule.
+The mean prediction $\widehat{f}_{\text{mean}}$ on the interval $t\in[0,1]$ is given by:
+[…]
+[…]
+The NLM compares the squared deviation between the prediction function and the linear secant to the squared deviation between the prediction function and the mean prediction:
+[…]
+Fig. 2 illustrates the setting for multivariate feature changes. The NLM can be approximated via numerical integration, e.g., via Simpson’s rule.
 Figure 1: Illustration by Scholbeck et al.
-(2022) of a univariate FME (blue bar) given the prediction function (yellow) and linear secant (green).
-The NLM indicates how well the secant can explain the prediction function - which is inversely proportional to the integral visualized in red - in relation to how well the most uninformative baseline model - the average prediction - can explain the prediction function.
+(2024) of a univariate FME (blue) given the prediction function (black) and linear secant (orange, dashed).
+The NLM indicates how well the secant can explain the prediction function (inversely proportional to the purple area) compared to how well the most uninformative baseline model (the average prediction) can explain the prediction function.
 Figure 2: Illustration of the multivariate NLM by Scholbeck et al.
@@
 Left: An exemplary bivariate prediction function and two points to compute an FME.
-Our observation is located at point (-5, -5), and we create the shortest path through the feature space to reach point (5, 5), which consists of directly proportional changes in both features.
-Above the path, we see the linear secant and the non-linear prediction function.
+Consider an observation $\bm{x}=(-5,-5)$ and step size vector $\bm{h}_{S}=(10,10)$.
+We create the shortest path through the feature space to reach the point (5, 5), which consists of directly proportional changes in both features.
+Above the path, we see the linear secant (orange, dashed) and the non-linear prediction function (black).
 Right: The multivariate change in feature values can be parameterized as a percentage $t$ of the step size $\bm{h}_{S}$.
-As a result, the deviation between prediction function and linear secant, as well as the deviation between the prediction function and mean prediction both correspond to a line integral which can be numerically approximated, for instance via Simpson’s rule.
-The NLM indicates how well the linear secant is able to explain the prediction function, compared to the baseline model of using the mean prediction.
+The deviation between the prediction function and the linear secant, as well as the deviation between the prediction function and mean prediction, both correspond to a line integral.
+The NLM indicates how well the linear secant can explain the prediction function, compared to the baseline model of using the mean prediction.
 A value of 1 indicates perfect linearity, where the linear secant is identical to the prediction function.
-For a value of 0, the mean prediction is able to explain the prediction function as equally well as the secant.
+For a value of 0, the mean prediction can explain the prediction function to the same degree as the secant.
 For negative values, the mean prediction better explains the prediction function than the linear secant (severe non-linearity).
@@
 Conditional average marginal effect
-In order to receive a global model explanation akin to a beta coefficient in linear models, local FMEs can be averaged to a global average marginal effect (AME).
-However, this does not accurately represent the nuances of a non-linear prediction function.
-Given a data set $\mathcal{D}$, the conditional average marginal effect (cAME) estimator applies to a subgroup of observations $\mathcal{D}_{[\;]}$:
+To receive a global model explanation akin to a beta coefficient in linear models, local FMEs can be averaged to the AME.
+Mehrabi et al.
+(2021) define an aggregation bias as drawing false conclusions about individuals from observing the entire population.
+Given a data set $\mathcal{D}$, the conditional average marginal effect (cAME) estimator applies to a subgroup of $n_{[\;]}$ observations, denoted by $\mathcal{D}_{[\;]}$:
 $\text{cAME}_{\mathcal{D}_{[\;]},\bm{h}_{S}}$ $=\widehat{\mathbb{E}_{\bm{X}_{[\;]}}\left[\text{FME}_{\bm{X}_{[\;]},\bm{h}_{S}}\right]}$
-$=\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{f}\left(\bm{x_{S}}^{(i)}+\bm{h}_{S},\bm{x_{-S}}^{(i)}\right)-\widehat{f}\left(\bm{x^{(i)}}\right)\right)\,\forall i:\bm{x^{(i)}}\in\mathcal{D}_{[\;]}$
-Although this estimator can be applied to arbitrary subgroups, an interpretation task demands searching for subgroups with homogeneous local model explanations.
-Desiderata for such subgroups include within-group effect homogeneity, between-group effect heterogeneity, full segmentation, non-congruence, confidence, and stability.
-A viable option to partition $\mathcal{D}$ is to run a recursive partitioning algorithm on $\mathcal{D}$ with FMEs as the target which aims at finding subgroups where FMEs are similar.
+$=\frac{1}{n_{[\;]}}\sum_{i\,:\,\bm{x}^{(i)}\in\mathcal{D}_{[\;]}}\left[\widehat{f}\left(\bm{x}_{S}^{(i)}+\bm{h}_{S},\bm{x}_{-S}^{(i)}\right)-\widehat{f}\left(\bm{x}^{(i)}\right)\right]$ (2)
+Although this estimator can be applied to arbitrary subgroups, we aim to find subgroups with cAMEs that counteract the aggregation bias.
+Desiderata for such subgroups include within-group effect homogeneity, between-group effect heterogeneity, full segmentation, non-congruence, confidence, and stability (Scholbeck et al., 2024).
+In other words, we aim to partition the data into subgroups that explain variability in the FMEs.
+A viable option to partition $\mathcal{D}$ is to run RP on $\mathcal{D}$ with FMEs as the target.
 For instance, in fmeffects, both rpart (Therneau and Atkinson, 2019) and ctree() from partykit (Hothorn and Zeileis, 2015) are supported to find subgroups.
+Related work
+Model-agnostic interpretations
+The basic mechanism behind model-agnostic methods is to probe the model with different feature values, a methodology similar to a model sensitivity analysis (Scholbeck et al., 2020, 2023).
+The basis of explaining models is to determine the direction and magnitude of the effect of features on the predicted outcome (Casalicchio et al., 2019; Scholbeck et al., 2020, 2024).
+The individual conditional expectation (ICE) (Goldstein et al., 2015), partial dependence (PD) (Friedman, 2001), accumulated local effects (ALE) (Apley and Zhu, 2020), Shapley values (Štrumbelj and Kononenko, 2010; Lundberg and Lee, 2017; Covert et al., 2020) and local interpretable model-agnostic explanations (LIME) (Ribeiro et al., 2016) are some of the most popular model-agnostic explanation methods for ML models.
+Notably, counterfactual explanations (Wachter et al., 2018) represent the reverse of the FME, indicating the smallest necessary change in feature values to reach a targeted prediction.
+FMEs complement the literature by allowing for a unique combination of local, regional, and global model explanations.
+Furthermore, while most methods (including the ICE, PD, ALE, or Shapley values) provide explanations in terms of prediction levels, FMEs provide explanations in terms of prediction changes.
+LIME is based on training a local and interpretable surrogate model whose coefficients can also provide an interpretation in terms of prediction changes.
+Scholbeck et al.
+(2024) highlighted differences between both approaches: notably, while surrogate models introduce additional uncertainty connected with the estimation of the surrogate, FMEs are motivated by the goal of stable and comprehensible model insight.
+Furthermore, locally estimated FMEs can be aggregated within subgroups and entire data sets for regional and global explanations.
+Around the same time, regional aggregations have also been introduced for ICE curves, for example (Britton, 2019; Herbinger et al., 2022; Molnar et al., 2024).
+Relationship between individual conditional expectation and forward marginal effect
+Scholbeck et al.
+(2024) illustrated a relationship between the ICE / PD and the FME / AME.
+In general, the FME can be seen as the difference between two locations on an ICE.
+The AME corresponds to the difference between two locations on the PD only for a function that is linear in the feature of interest.
+Therefore, the following relationship between the ICE and FME is worth noting here.
+The ICE can be considered a one-way sensitivity function that indicates the effects of varying a set of features indexed by $S$ while the remaining ones are kept constant:
+$\text{ICE}_{\bm{x},S}(\bm{x}^{\ast}_{S})=\widehat{f}(\bm{x}^{\ast}_{S},\bm{x}_{-S})$
+For an instance $\bm{x}$, the prediction after increasing $\bm{x}_{S}$ by $\bm{h}_{S}$ is also a value of the ICE:
+$\text{FME}_{\bm{x},\bm{h}_{S}}$ $=\widehat{f}(\bm{x}_{S}+\bm{h}_{S},\bm{x}_{-S})-\widehat{f}(\bm{x})$
+$=\text{ICE}_{\bm{x},S}(\bm{x}_{S}+\bm{h}_{S})-\text{ICE}_{\bm{x},S}(\bm{x}_{S})$
+Related work on marginal effects
+MEs have a long history in applied statistics and the Stata programming language (StataCorp, 2023).
+Initially implemented by Bartus (2005), the margins() command is now fully integrated into Stata and provides comprehensive capabilities for various computations and visualizations of statistical models such as (generalized) linear models (Williams, 2012).
+MEs are typically defined in terms of derivatives of the model w.r.t.
+a feature.
+For instance, this variant is the default approach to interpret models in econometrics (Greene, 2019).
+The FME is the less commonly used definition (Scholbeck et al., 2024; Mize et al., 2019).
+Note that—in contrast to forward differences—derivatives are not suitable to explain piecewise constant prediction functions such as tree-based models.
+In recent years, MEs have gained traction in the R community.
+The R package margins (Leeper, 2018) was the first port of Stata’s margins() command to R.
+Other packages related to MEs include ggeffects (Lüdecke, 2018) and marginaleffects (Arel-Bundock, 2023).
+In particular, marginaleffects can also return FMEs (although under different terminology).
+Our package, fmeffects, mainly differs from marginaleffects in two aspects:
+Implementing new theory surrounding FMEs:
+The fmeffects package is the first software implementation of the theory surrounding model-agnostic FMEs as introduced by Scholbeck et al.
+(2024).
+Although packages such as marginaleffects support the computation of FMEs and other quantities, fmeffects is specifically designed for FMEs with unique features such as implementations of the NLM, the cAME via RP, and novel visualization methods.
+Model-agnostic black box interpretations:
+It follows that fmeffects is targeted at model-agnostic explanations of non-linear or intransparent models.
+Whereas existing theory on MEs (and packages such as marginaleffects) focuses on classical statistical modeling in combination with statistical inference (see, for instance, Breiman (2001) comparing statistical modeling culture with ML), FMEs (and thus fmeffects) are comparable to methods and software from the literature on interpretable ML such as the ICE, PD, ALE, or LIME.
+This does not imply that marginaleffects cannot be used for black box interpretations.
+As mentioned in the previous point, it also supports the computation of FMEs, e.g., in combination with mlr3, but the focus of fmeffects lies on the interpretation of black box models through a specialized and targeted range of novel capabilities.
+Advantages and limitations of forward marginal effects
+Advantages
+Although the ICE and the FME are closely related, the latter provides several novel ways to generate insights into the model:
+  • Univariate changes in feature values: FMEs are comparable to ICE curves for univariate changes in feature values.
+In certain scenarios, however, they may provide more comprehensible visualizations of effects for individual instances (see Fig.
+4 for an example).
+  • Bivariate changes in feature values: The ICE and PD also provide insight into the sensitivity of the model prediction for variations in two features, which is visualized as a heatmap (see Fig.
+7).
+However, it is difficult to visually compare the ICE of many different observations (which correspond to heatmaps as well).
+Although the ICE provides insight into a larger variation in feature values, while the FME only considers a single tuple of changes in feature values, bivariate FMEs can be easily compared visually (see Fig.
+6).
+  • Higher-order changes in feature values: If we evaluate the sensitivity of the prediction for changes in more than two feature values, virtually every visualization method breaks down.
+In this case, FMEs still provide comprehensible model explanations that can be aggregated in various ways (see Fig.
+10).
+  • Local fidelity assessment: The locally restricted change in feature values for the FME facilitates evaluations of the fidelity of the model explanation (e.g., via the NLM).
+In other words, the NLM allows us to describe how well the FME summarizes the local shape of the prediction function in a single value.
+See Fig.
+8 for a visualization of NLM values for different observations.
+  • Comprehensible regional explanations: Although regional explanations have been first proposed in the context of grouping ICE curves (Herbinger et al., 2022; Britton, 2019), they more easily apply to scalar model explanations such as FMEs.
+Essentially, a regional model explanation represents a group of observations or a subspace of the feature space where model explanations are relatively homogeneous.
+Such groupings are easily achievable via RP or other techniques that do not require functional target values such as ICEs.
+  • Avoiding extrapolation: The ICE is computed on the entire feature range (see, e.g., Fig.
+4), which is likely to result in model extrapolations.
+By its nature, the FME is typically used with small step sizes relative to the feature range, which naturally avoids model extrapolations.
+Limitations
+  • Step size selection: The step size fundamentally influences effects and the model interpretation.
+Although FMEs for different step sizes can be computed and visualized in an exploratory manner, some level of prior reasoning about what step sizes to use is recommended.
+  • Decision tree instability for cAME: Although not a shortcoming of the FME itself, subgroups found by RP to compute cAMEs are subject to a high variance.
+This may be counteracted by stabilizing the split search, e.g., by considering statistical significance of tree splits or resorting to different algorithms to find subgroups.
+  • Non-linearity assessment for proportional feature changes: For multi-dimensional feature changes, the NLM only considers equally proportional changes in all features.
+On causal interpretations and avoiding model extrapolations
+Note that model-agnostic techniques, including FMEs, explain associations between the target and the features within the model.
+In the absence of additional assumptions, such associations cannot be interpreted as causes and effects (Molnar et al., 2022).
+For instance, increasing the value of a feature $x_{1}$ may always be accompanied by an increase in the target, but it may be the target $y$ that causes $x_{1}$ to increase.
+Another typical scenario is the presence of confounding factors that influence both $y$ and $x_{1}$.
+Finally, $x_{1}$ may only (or also) influence a mediator $x_{2}$, which in turn influences $y$.
+This does not, however, make model interpretations obsolete.
+More importantly, as highlighted by Adadi and Berrada (2018), model interpretations can be used to gain knowledge, debug, audit, or justify the model and its predictions.
+Throughout this paper, we will model the effects of environmental influences on the number of daily bike rentals in Washington, D.C.
+For our estimated model, a drop in humidity by 10 percentage points has a considerable effect on the predicted number of daily bike rentals (see Fig.
+5).
+This effect cannot be assumed to be causal, as humidity is physically influenced by the outside temperature, which will also affect people’s choice to rent a bike.
+Here, temperature is a confounder that influences both humidity and daily bike rentals.
+However, the business renting out bikes can still use the associations found by a model with a good predictive performance to control the optimal number of bikes at their disposal.
+This is conditional on the model’s ability to accurately predict the target for the given feature vector, requiring us to avoid model extrapolations, which correspond to predictions within areas of the feature space where the model has not seen much or any training data.
+This issue is closely linked to the multivariate distribution of the training data; in our example, a change in humidity is likely to be accompanied by a change in temperature as well, which we somewhat circumvent (depending on the magnitude of the step size) when making isolated changes to humidity.
+One may disregard this issue and deliberately predict in areas of the feature space the model has not seen during training.
+The resulting FMEs will still be valid model descriptions but, as explained above, they are likely to be bad descriptions of the data generating process.
+Model extrapolations negatively impact many model-agnostic interpretation methods (Hooker, 2004b, a, 2007; Hooker et al., 2021; Molnar et al., 2022).
+For example, Apley and Zhu (2020) demonstrated how PD plots suffer from extrapolation issues and introduced ALE plots as a solution to this problem.
+Scholbeck et al.
+(2024) illustrated the perils of model extrapolations for FMEs specifically and discussed possible options.
+One option in particular is also implemented in fmeffects: points outside the multivariate envelope (meaning the Cartesian product of all observed feature ranges) of the training data can be excluded from the analysis.
+This directly relates to the selection of small step sizes relative to the feature range, as large step sizes will result in a point falling outside the envelope.
+When using extrapolation prevention methods, note that we consider different sets of points for different step sizes, which differs from the usage of MEs in other contexts (see, for instance, the package marginaleffects for a comparison).
+The exclusion of points only impacts aggregations of FMEs, i.e., the cAME and AME.
+As discussed in the section on Forward marginal effects, this also affects the computation of categorical AMEs.
+In Eq.
+(1) and Eq.
+(2), the AME and cAME are formulated as estimators of the expected global or regional (concerning a subspace) effects.
+The fewer observations we are considering for an average, the larger the variance of the estimate.
 User interface and package handling
 Local explanations
-The fme() function is the central user interface, which - apart from certain parameters - only requires a pre-trained model and a data set (see section Design and options for extensions for details).
-Further arguments specify the target, feature(s), step size(s), whether to compute NLM values for each FME, and specification of an extrapolation detection method [^3] Model extrapolations are predictions in areas with a low density of training observations.
-Scholbeck et al.
-(2022) provide a short introduction to this topic..
-The fme() function initiates construction of and computations for a ForwardMarginalEffects object without requiring the user to know R6 (Chang, 2021) syntax.
+The fme() function is the central user interface.
+It mainly requires a pre-trained model and a data set (see section Design and options for extensions for details).
+Further control parameters include a list of features and step sizes, whether to compute NLM values for each FME, and an extrapolation detection method.
+The fme() function initiates the construction and computations of a ForwardMarginalEffect object without requiring the user to know R6 (Chang, 2021) syntax.
 For this use case, we train a random forest from the ranger package (Wright and Ziegler, 2017) on the bike sharing data set (Fanaee-T, 2013) using mlr3.
-Note that models trained via tidymodels and caret are also supported.
-Our goal is to predict and explain the daily bike rental demand in Washington D.C.
-based on features such as the outside temperature, wind speed, or humidity.
+Note that models trained via tidymodels and caret are also supported, as well as models trained via lm(), glm(), and gam().
+We aim to predict and explain the daily bike rental demand in Washington, D.C., based on features such as the outside temperature, wind speed, or humidity.
 We first train the model:
@@
 Then, we simply pass the trained model, evaluation data, and remaining parameters to the fme() function.
-It returns a ForwardMarginalEffects object which can be analyzed via summary() and visualized via plot() (see Fig.
+It returns a ForwardMarginalEffect object, which can be analyzed via summary() and visualized via plot() (see Fig.
 3).
 Here, the outside temperature is raised by 5 degrees Celsius ceteris paribus.
-Each dot represents a single instance.
-The arrow above the horizontal axis indicates the step size and direction of the feature change in relation to the feature range.
-The horizontal value of each instance indicates its observed feature value in the provided data set; the vertical axis value is the corresponding FME.
-Let us single out the observation with the largest associated FME (brown colored and enlarged in Fig.
-3).
-This observation corresponds to a single day with a recorded temperature of 13.12 degrees Celsius.
-Increasing the temperature by 5 degrees Celsius on this particular day results in an additional 125 predicted bike rentals.
+To avoid overplotting values, each hexagon represents a local average of FMEs.
+Users can easily access the data used by all plot functions to implement their own visualizations.
+Let us single out the observation with the largest associated FME.
+This observation corresponds to a single day with a recorded temperature of 8 degrees Celsius.
+Increasing the temperature by 5 degrees Celsius on this particular day results in 2699 additional predicted bike rentals.
 We plot such model explanations for the entire data set and average FMEs to receive a global model explanation.
-The AME - the global average of FMEs - is 14.5: An increase in temperature by 5 degrees Celsius results in an average change in predicted daily bike rentals by 14.5.
-> effects = fme(+   model = forest,+   data = bikes,+   target = "count",+   feature = "temp",+   step.size = 5,+   ep.method = "envelope")> summary(effects)
-Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5Extrapolation point detection:  envelope, EPs: 81 of 727 obs. (11 %)Average Marginal Effect (AME):  14.4844
-> plot(effects)
+The AME—the global average of FMEs—is 307: an increase in temperature by 5 degrees Celsius results in an average increase of 307 predicted daily bike rentals.
+> effects.univariate.temp = fme(+   model = forest,+   data = bikes,+   features = list("temp" = 5),+   ep.method = "envelope")> summary(effects.univariate.temp)
+Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5Extrapolation point detection:  envelope, EPs: 48 of 731 obs.
+(7 %)Average Marginal Effect (AME):  307.3275> plot(effects.univariate.temp)
 Figure 3: Plot of univariate FMEs for feature ‘temp’ and step size 5.
-Each dot represents an observation.
-The horizontal value represents the observed value of feature ‘temp’.
-We move each observation’s ‘temp’ value according to the arrow’s direction and length (in this case, 5 units to the right).
-The respective FME value is indicated by the vertical value of each point.
-The horizontal bar indicates the AME (the global FME average).
-An additional smoothing function facilitates interpretations by modeling an approximate pattern of FMEs across the feature range.
-We single out one observation (brown colored and enlarged) to demonstrate a local interpretation.
-Let us take a moment to compare the FME plot with the individual conditional expectation (ICE) (Goldstein et al., 2015) and partial dependence (PD) (Friedman, 2001) plot generated by the R package iml (Molnar et al., 2018).
-The ICE and PD are one of the most popular and established model-agnostic ways to interpret predictive models (Molnar, 2022).
-The ICE is a local model explanation and represents the prediction for an observation where only the features of interest are varied (in this case only ‘temp’).
+Each hexagon represents a local FME average.
+The horizontal value represents the observed feature value of ‘temp’.
+Each observation’s ‘temp’ value is moved according to the arrow’s direction and length.
+The vertical value of each hexagon indicates the FME value associated with that feature change.
+The horizontal bar indicates the AME.
+The shade of the hexagon implies how many observations it contains.
+A smoothing function facilitates interpretations by modeling an approximate pattern of FMEs across the feature range.
+Let us take a moment to compare the FME plot with the combined ICE and PD plot generated by the R package iml (Molnar et al., 2018) (see Fig.
+4).
+This is one of the most popular and established model-agnostic ways to interpret predictive models (Molnar, 2022).
+The ICE is a local model explanation and represents the prediction for an observation where only the features of interest are varied (in this case, only ‘temp’).
 The PD is the average of ICEs (in the univariate case, the vertical average) and indicates the global, average prediction when a subset of features is varied for all observations.
-From Fig.
-4, we gather that in contrast to FMEs, the ICE and PD plot is more difficult to interpret.
 Although we can see a rough trajectory of the feature influence on local and average predictions, it is difficult to pinpoint the exact effects of changing ‘temp’ on the prediction for single observations.
-Furthermore, ICE curves are more likely to be subject to model extrapolations, a result from predicting in areas where the model was not trained on a sufficient amount of data.
+Furthermore, ICE curves are more likely to be subject to model extrapolations, a result of predicting in areas where the model was not trained on a sufficient amount of data.
 Figure 4: An ICE and PD plot for feature ‘temp’ generated by the R package iml.
-Each black curve (an ICE) represents predictions for a single instance while only feature ‘temp’ varies.
-The yellow curve (the PD) is the vertical average of ICEs and represents the average, isolated influence of ‘temp’ on predictions.
-As opposed to FMEs, this plot is substantially more difficult to interpret.
+Each solid blue curve (an ICE) represents predictions for a single instance while only ‘temp’ varies.
+The dashed black curve (the PD) is the vertical average of ICEs and represents the average, isolated influence of ‘temp’.
+FMEs allow for positive or negative step sizes.
+For instance, let us investigate the effects of an isolated drop in humidity by 10 percentage points.
+We can observe an AME of 108 additional predicted bike rentals a day.
+Individual effects tend to be larger the higher the humidity on that particular day.
+> effects.univariate.humidity = fme(+   model = forest,+   data = bikes,+   features = list("humidity" = -0.1),+   ep.method = "envelope")> summary(effects.univariate.humidity)
+Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  humidity, -0.1Extrapolation point detection:  envelope, EPs: 1 of 731 obs. (0 %)Average Marginal Effect (AME):  108.0477
+> plot(effects.univariate.humidity)
+Figure 5: Univariate FMEs for a drop in humidity by 10 percentage points. Especially for high humidity values, the drop results in a considerable increase in predicted daily bike rentals.
 In many applications, we are interested in interactions of features on the prediction.
-We diagnosed a univariate effect of ‘temp’ on the predicted amount of bike rentals, but effects may be enhanced or mitigated when other features are varied at the same time.
-For instance, humidity has a considerable impact on the effects of heat (Russo et al., 2017).
+Until now, we only analyzed the univariate effects of ‘temp’ and ‘humidity’ on the predicted amount of bike rentals.
+However, potential interactions between features may exist.
 We evaluate an increase in temperature by 5 degrees Celsius and a simultaneous drop in humidity by 10 percentage points (see Fig.
-5).
+6).
 For a bivariate change in feature values, the two arrows depict the direction and magnitude of the feature change in the respective variable.
-Each point is located at the observation’s observed feature values in the provided data set.
+As in the univariate case, we plot local averages within hexagons to avoid overplotting values.
+The location of the hexagon is determined by the observations’ observed feature values in the provided data set.
 Its color indicates the FME associated with the bivariate feature change.
-An increase in the outside temperature by 5 degrees Celsius and a drop in humidity by 10 percentage points is associated with an AME of 19.14.
-The isolated univariate effect of increasing the outside temperature is boosted by the simultaneous drop in humidity.
-> effects.bivariate = fme(model = forest,+               data = bikes,+              target = "count",+              feature = c("temp", "humidity"),+              step.size = c(5, -0.1),+              ep.method = "envelope")> summary(effects.bivariate)
-Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5  humidity, -0.1Extrapolation point detection:  envelope, EPs: 82 of 727 obs.
-(11 %)Average Marginal Effect (AME):  19.1418
+An increase in the outside temperature by 5 degrees Celsius and a simultaneous drop in humidity by 10 percentage points is associated with an AME of 414.
+The univariate AMEs roughly add up to the bivariate AME, indicating that, on average, there is no additional interaction between both feature changes on the prediction.
+> effects.bivariate = fme(+   model = forest,+   data = bikes,+   features = list("temp" = 5, "humidity" = -0.1),+   ep.method = "envelope")> summary(effects.bivariate)
+Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5  humidity, -0.1Extrapolation point detection:  envelope, EPs: 49 of 731 obs.
+(7 %)Average Marginal Effect (AME):  413.6163
 > plot(effects.bivariate)
@@
 FMEs are highly heterogeneous.
-We can see mostly positive effects, especially for observations with combinations of lower ‘temp’ and ‘humidity’ values.
-For various really warm days, the model even predicts a negative effect.
+We can see mostly positive effects, especially for observations with combinations of medium ‘temp’ and ‘humidity’ values.
 Let us repeat the same procedure as for univariate feature changes and compare the FME plot to an alternative option, the bivariate PD plot (see Fig.
-6).
-As opposed to the novel visualization with FMEs, the PD plot only visualizes the average, global effect of changing both features on the predicted amount of bike rentals, and it does not inform about the distribution of observed feature values, thus not allowing us to evaluate effects of increasing one feature and decreasing another at the same time.
-Furthermore, the PD plot - like its univariate counterpart - is more likely to be subject to model extrapolations due to predicting for many feature value combinations the model has not seen during training.
-Figure 6: A bivariate PD plot (created via the R package iml), one of the few available visualization tools for interactions.
-We can see the global interaction between ‘temp’ and ‘humidity’ on the predicted amount of bike rentals.
-Plugging in larger values for both features ceteris paribus results in more predicted bike rentals on average.
-As opposed to bivariate FMEs, we cannot investigate multiple local effects, nor can we see the actual distribution of observed feature values, and as a result, we cannot evaluate the effects of increasing one feature and decreasing another at the same time.
-Let us now proceed to investigations of non-linearity.
-Although some sense of non-linearity can be derived from ICE curves (see Fig.
-4), it is hard to quantify, and summarizing non-linearity for an ICE with a large variation in the feature of interest would be somewhat meaningless.
+7).
+As opposed to the novel visualization with FMEs, the PD plot only visualizes the average, global effect of changing both features on the predicted amount of bike rentals.
+It does not inform us about the distribution of observed feature values, thus not allowing us to evaluate the effects of increasing one feature and decreasing another simultaneously.
+Figure 7: A bivariate PD plot (created via the R package iml), visualizing the global interaction between ‘temp’ and ‘humidity’ on the predicted amount of bike rentals.
+Plugging in medium to large values for ‘temp’ and low to medium values for ‘humidity’, ceteris paribus, results in more predicted bike rentals on average.
+As opposed to bivariate FMEs, we cannot investigate multiple local effects, nor can we see the actual distribution of observed feature values.
+As a result, we cannot evaluate the effects of increasing one feature and decreasing another simultaneously.
+Let us now proceed to investigate non-linearity.
+Non-linearity can be visually assessed for ICE curves (see Fig.
+4), but it is hard to quantify and would be somewhat meaningless for a large variation in the feature of interest.
 Furthermore, for bivariate or higher-dimensional changes in feature values, we lose any option for visual diagnoses of non-linearity.
-In contrast, the NLM can be computed for FMEs with continuous step sizes, regardless of their dimensionality:
-> effects.bivariate.nlm = fme(model = forest,+                       data = bikes,+                       target = "count",+                       feature = c("temp", "humidity"),+                       step.size = c(5, -0.1),+                       ep.method = "envelope",+                       compute.nlm = TRUE)> plot(effects.bivariate.nlm, with.nlm = TRUE)
-Figure 7: Adding NLM computations to the FME plot.
-Each dot in the left and right plot represents a single observation and its respective FME and NLM value.
-We can see both linear (bright, yellow colored NLM) and non-linear FMEs (dark, blue colored NLM).
+In contrast, the NLM can be computed for FMEs with continuous step sizes, regardless of dimensionality.
+The average non-linearity measure (ANLM) is 0.36, indicating that the linear secant, on average, is a bad descriptor of the FME.
+> effects.bivariate.nlm = fme(+   model = forest,+   data = bikes,+   features = list("temp" = 5, "humidity" = -0.1),+   ep.method = "envelope",+   compute.nlm = TRUE)> effects.bivariate.nlmForward Marginal Effects ObjectFeatures & step lengths:  temp, 5  humidity, -0.1Average Marginal Effect (AME):  413.6163Average Non-Linearity Measure (ANLM):  0.36> plot(effects.bivariate.nlm, with.nlm = TRUE)
+Figure 8: Adding NLM computations to the FME plot. Each hexagon in the left and right plots represents a local average of FME and NLM values, respectively.
 Fig.
-7 simply contrasts each observation’s NLM value with its FME value.
-In this case, we can see both non-linear FMEs (darker, blue colored NLM values) and linear FMEs (brighter, yellow colored NLM values).
-This is not surprising, considering that the bike sharing data consists of many features, each being capable to influence the effect of changes in ‘temp’ and ‘humidity’ on the prediction.
+8 simply contrasts FME values with the corresponding NLM values.
+In this case, we can see both non-linear FMEs (whiter NLM) and linear FMEs (deep blue-colored NLM).
+We could now, for instance, focus on interpreting linear FMEs.
+All FMEs depicted in Fig.
+9 have an NLM of 0.9 or higher, meaning that they almost fully describe the model prediction for proportional changes in ‘temp’ and ‘humidity’.
+Figure 9: Visualizing FMEs with an NLM $\geq$ 0.9.
+An advantage of FMEs is their ability to provide comprehensible model insight even when exploring higher-order feature changes.
+Let us factor in a third feature change, now simultaneously reducing windspeed by 5 miles per hour, and visualize the distribution of FME and NLM values.
+We can see that in addition to an increase in temperature and a decrease in humidity, a decrease in windspeed further boosts the average number of predicted daily bike rentals.
+> effects.trivariate.nlm = fme(+   model = forest,+   data = bikes,+   features = list("temp" = 5, "humidity" = -0.1, "windspeed" = -5),+   ep.method = "envelope",+   compute.nlm = TRUE)> summary(effects.trivariate.nlm)Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5  humidity, -0.1  windspeed, -5Extrapolation point detection:  envelope, EPs: 117 of 731 obs.
+(16 %)Average Marginal Effect (AME):  537.7385Average Non-Linearity Measure (ANLM):  0.33> plot(effects.trivariate.nlm, with.nlm = TRUE)
+Figure 10: Adding a third feature change, a drop in windspeed by 5 miles per hour, and visualizing the distribution of FME and NLM values.
+For the NLM plot, negative NLMs are binned as 0.
+It follows that the ANLM value in the plot differs from the raw ANLM in the summary output.
 So far, we have only evaluated changes in continuous features.
-In many applications, we are often concerned with switching categories of categorical features, a way of counterfactual thinking inherent to the human thought process.
-Note that in spite of the allure of switching categories of categorical features, one needs to be aware of potential model extrapolations.
+In many applications, we are concerned with switching categories of categorical features, a way of counterfactual thinking inherent to the human thought process.
+Note that despite the allure of switching categories of categorical features, one needs to be aware of potential model extrapolations.
 To illustrate this, we switch each non-rainy day’s precipitation status to rainfall.
-Rainfall has an average isolated effect of lowering daily rentals by 57 bikes (see Fig.
-8).
-> effects.categ = fme(model = forest,+                data = bikes,+                target = "count",+                feature = "weather",+                step.size = "rain")> summary(effects.categ)Forward Marginal Effects ObjectStep type:  categoricalFeature & reference category:  weather, rainExtrapolation point detection:  none, EPs: 0 of 657 obs.
-(0 %)Average Marginal Effect (AME):  -57.3322> plot(effects.categ)
-Figure 8: Distribution of categorical FMEs resulting from switching each non-rainy day’s precipitation status to rain. On average, rainfall lowers predicted bike rentals by 57 bikes per day.
+Rainfall has an average isolated effect of lowering daily rentals by 803 bikes (see Fig.
+11).
+> effects.categ = fme(+   model = forest,+   data = bikes,+   features = list("weather" = "rain"))> summary(effects.categ)Forward Marginal Effects ObjectStep type:  categoricalFeature & reference category:  weather, rainExtrapolation point detection:  none, EPs: 0 of 710 obs.
+(0 %)Average Marginal Effect (AME):  -802.8716> plot(effects.categ)
+Figure 11: Distribution of categorical FMEs resulting from switching each non-rainy day’s precipitation status to rain. On average, rainfall lowers predicted bike rentals by 803 bikes per day.
 Regional explanations
@@
 This is where regional explanations come into play.
-We aim to find subgroups with more homogeneous FME values, thereby describing the behavior of the model not in terms of a global average, but in terms of multiple regional averages (cAMEs).
-In fmeffects, this can be achieved by further processing the ForwardMarginalEffects object containing FMEs (and optionally NLM values) using the came() function.
+We aim to find subgroups with more homogeneous FME values, thereby describing the behavior of the model not in terms of a global average but in terms of multiple regional averages (cAMEs).
+In fmeffects, this can be achieved by further processing the ForwardMarginalEffect object containing FMEs (and optionally NLM values) using the came() function.
 This returns a Partitioning object (in this case, an object of the class "PartitioningCTREE", a subclass of the abstract class "Partitioning", see later section on Design and options for extensions).
-For the univariate change in temperature by 5 degrees Celsius, we decide to search for exactly 3 subgroups [^4] This value is to be set by the user depending on how many regional explanations are to be found.
+For the univariate change in temperature by 5 degrees Celsius, we decide to search for precisely 2 subgroups [^2] This value is to be set by the user depending on how many regional explanations are to be found.
+Alternatively, we can search for a pre-defined SD of FMEs inside the terminal nodes.
 How many subgroups can be found depends on the data and predictive model.
@@
 A summary of the created object informs us about the number of observations, cAME, and standard deviation (SD) of FMEs inside the root node and leaf nodes (the found subgroups).
-We succeed in finding subgroups with lower SDs while maintaining an appropriate sample size.
-The root node SD of 22.5 can be successfully split down to 21.5, 8.5, and 17.7 inside the subgroups.
+We succeeded in finding subgroups with lower SDs while maintaining an appropriate sample size.
+The root node SD of 611 can be successfully split down to 437 and 355 within the subgroups.
 By visualizing the tree, we can see how the data was partitioned.
-Lower temperatures are associated with larger increases in predicted bike rentals when the outside temperature is raised.
-For outside temperatures lower than 17 degrees Celsius and working days, we can observe the largest cAME of 32 additional bike rentals per day and a cAME of 11 additional daily bike rentals for weekends.
-For outside temperatures above 17 degrees Celsius, there is only a slight effect of 1.7 additional bikes per day:
-> subspaces = came(effects = effects, number.partitions = 3)> summary(subspaces)
-PartitioningCtree of an FME objectMethod:  partitions = 3   n      cAME   SD(FME) 646 14.484353 22.499156 * 230 32.265791 21.546565 130 11.172507  8.554322 286  1.689979 17.749573---* root node (non-partitioned)AME (Global): 14.4844
+For cooler outside temperatures equal to or lower than $\approx$ 16 degrees Celsius, we can observe a positive cAME of 728 additional bike rentals per day.
+On warmer days with a temperature above $\approx$ 16 degrees Celsius, the model predicts 196 less bike rentals a day when the outside temperature increases by 5 degrees.
+> subspaces = came(effects = effects.univariate.temp, number.partitions = 2)> summary(subspaces)
+PartitioningCtree of an FME objectMethod:  partitions = 2   n      cAME  SD(fME) 683  307.3275 611.0778 * 372  728.3942 437.0463 311 -196.3278 354.5090---* root node (non-partitioned)AME (Global): 307.3275
 > plot(subspaces)
-Figure 9: Finding subgroups of observations with a decision tree. Each leaf node visualizes one subgroup, the number of observations, the cAME, and the SD of FMEs indicating FME heterogeneity.
+Figure 12: Using a decision tree to find subgroups of observations with more homogeneous FMEs of increasing ‘temp’ by 5 degrees Celsius.
+Each leaf node visualizes one subgroup, the number of observations, the cAME, and the SD of FMEs indicating FME homogeneity.
 Global explanations
-When to search for regional explanations thus depends on the heterogeneity of local effects, which can be summarized for the entire model via the ame() function.
+When to search for regional explanations thus depends on the heterogeneity of local effects.
+The ame() function provides an appropriate summary for the entire model.
 It uses a default step size of 1 or 0.01 for small feature ranges.
 For categorical FMEs, it uses every observed category as a reference category.
-The summary() function prints a compact model summary of each feature, a default step size, the AME, the SD of FMEs, 25% and 75% quantiles of FMEs, as well as the number of observations left after excluding extrapolation points.
-A large dispersion indicates heterogeneity of FMEs and thus, a small fidelity of the AME and possible benefits from searching for subgroups with varying cAMEs.
-A different workflow can, therefore, also consist of starting with the table generated by ame() and deciding which feature effects can be described by AMEs, which might be better describable by subgroups and cAMEs, and - if this has been without success - resorting to local model explanations.
+Alternatively, custom step sizes and subsets of features can be used.
+The summary() function prints a compact model summary of each feature, a default step size, the AME, the SD of FMEs, 25% and 75% quantiles of FMEs, as well as the number of observations left after excluding extrapolation points (EPs).
+A large dispersion indicates heterogeneity of FMEs and thus a small fidelity of the AME and possible benefits from searching for subgroups with varying cAMEs.
+A different workflow can, therefore, also consist of starting with the table generated by ame() and deciding which feature effects can be described by AMEs and which might be better describable by subgroups and cAMEs.
+If this has been unsuccessful, we can resort to local model explanations.
 Recall our example from the previous section on Regional explanations where we split FMEs associated with increasing temperature by 5 degrees Celsius.
-From the ame() summary, we gather that ‘temp’ has a relatively large SD in relation to its AME, and the interquartile range indicates a large spread of FMEs from -0.47 in the 25% quantile up to 4.74 in the 75% quantile, thus making it a promising candidate to find subgroups with more homogeneous FMEs.
-> ame.results = ame(model = forest, data = bikes, target = "count")> summary(ame.results)
-Model Summary Using Average Marginal Effects:      feature step size       AME      SD      0.25      0.75   n1      season    spring  -29.3415 30.6189  -38.2318   -6.3963 5482      season    summer    0.5742 21.4782   -8.3499   10.5607 5433      season      fall      9.97 27.4275   -1.9819    30.745 5394      season    winter   13.9861 23.6806    1.0266   23.6969 5515        year         0  -98.6521 67.2816 -156.4247    -19.63 3646        year         1   96.6155 60.1458     22.91  148.1915 3637       month         1    3.9954 13.4096   -1.2981    7.2207 7278     holiday     False   -1.5799 20.5126  -12.5858   14.0711  219     holiday      True  -12.0674 24.2779  -29.8895    6.4375 70610    weekday       Sat  -54.4458 49.1836  -85.8245  -15.5722 62211    weekday       Sun  -78.6803  51.462 -113.7695  -31.2021 62212    weekday       Mon   10.7612 28.3258   -8.7131   30.5779 62313    weekday       Tue   17.9565 24.7443    0.9637   33.5499 62514    weekday       Wed    20.721 22.6949    1.5863   34.3365 62315    weekday       Thu   19.7423 23.9554   -0.2864   34.7076 62416    weekday       Fri    1.6724 35.8582  -23.9194   29.7121 62317 workingday     False -204.2023 90.1427 -256.0286 -141.4579 49618 workingday      True  165.2412 64.8985  123.8688   213.351 23119    weather     clear   25.9084  42.223    3.6162   22.3112 28420    weather     misty    3.3301 33.2101   -9.0864    0.8842 51321    weather      rain  -57.3322 55.5617  -95.7276   -3.8571 65722       temp         1    2.3959  7.2201   -0.4717    4.7369 72723   humidity      0.01   -0.1681  2.4691   -0.2999    0.4329 72724  windspeed         1    0.0198  2.2861   -0.2156    0.2126 727
+From the ame() summary, we see that ‘temp’ has a relatively large SD in relation to its AME (here calculated with a step size of 1), and the interquartile range indicates a wide spread of FMEs from -20 in the 25% quantile up to 108 in the 75% quantile, which makes it a promising candidate to find subgroups with more homogeneous FMEs.
+> ame.results = ame(model = forest, data = bikes)> summary(ame.results)
+Model Summary Using Average Marginal Effects:      Feature step.size        AME       SD       0.25       0.75   n1      season    winter  -942.0906 466.3691 -1298.1011  -617.5663 5502      season    spring   136.2185 569.5307  -244.4237   650.0125 5473      season    summer   293.6264 549.2972   -42.7551   738.2056 5434      season      fall   533.5502 579.5541    52.3706  1138.0863 5535        year         0 -1899.4966 639.1695 -2354.1389 -1506.0582 3666        year         1  1790.6269 524.4711  1421.7925  2194.1396 3657     holiday        no     195.93  218.386   123.2468   228.0909  218     holiday       yes  -133.3134 154.8869  -201.3635   -25.1245 7109     weekday    Sunday   155.5219 188.8708     9.3486   252.0308 62610    weekday    Monday  -158.9218 215.5047  -263.2441    -4.8485 62611    weekday   Tuesday  -115.7316 193.4508  -197.7396    13.3208 62612    weekday Wednesday   -44.3056 173.8664  -115.5562    63.1344 62713    weekday  Thursday     16.005  161.125   -61.1673    89.5043 62714    weekday    Friday    57.1498 163.5602   -27.6519    128.752 62715    weekday  Saturday   103.7648 170.5678    -0.2044    178.493 62716 workingday        no   -42.8794 139.8572  -145.7104    66.2131 50017 workingday       yes    48.1298 158.3666   -60.2448   145.5003 23118    weather     misty  -221.5664 328.3458  -413.4363   -69.4238 48419    weather     clear   385.8674 347.6119   162.2048   476.8631 26820    weather      rain  -802.8716 384.2624 -1054.7158  -543.2614 71021       temp         1    58.0487 164.8714   -20.0019   108.4669 73122   humidity      0.01     -19.86  62.1753   -36.5407    10.4535 73123  windspeed         1   -24.7315  77.1757   -56.9247    13.7468 731
 Design and options for extensions
@@
 13 provides a visual overview of the core design.
-Major emphasis is given to the strategy and adapter design patterns (Gamma et al., 1994).
-Simply put, the strategy pattern decouples source code for algorithm selection at runtime into separate classes.
+The greatest emphasis is placed on the strategy and adapter design patterns (Gamma et al., 1994).
+Simply put, the strategy pattern decouples the source code for algorithm selection at runtime into separate classes.
 We repeatedly implement this pattern throughout the package by creating abstract classes whose subclasses implement specific functionalities.
-The adapter design pattern (also referred to as a “wrapper”) creates an interface for communication between two classes.
+The adapter design pattern (also called a “wrapper”) creates an interface for communication between two classes.
 Figure 13: Design overview of the fmeffects package, including methods that implement the main functionality of each class.
@@
 This allows users of fmeffects to use numerous predictive models such as random forests, gradient boosting, support vector machines, or neural networks.
+"PredictorLM" stores models returned by lm(), glm(), or gam().
 The package can be extended with novel model types by implementing a new subclass that stores the model, data, target, and is able to return predictions.
-  • "AverageMarginalEffects": A class to compute AMEs for each feature in the data (or a subset thereof).
-Internally, a new "ForwardMarginalEffects" object is used to compute FMEs and aggregate them.
+  • "AverageMarginalEffects": A class to compute AMEs for each feature in the data (or a subset of features).
+Internally, a new "ForwardMarginalEffect" object is used to compute and aggregate FMEs.
 For convenience, we implement a wrapper function ame() to facilitate object creation and to initiate computations without requiring user input in the form of R6 syntax.
-  • "ForwardMarginalEffects": The centerpiece class of the package.
-It keeps access to a Predictor, stores important information to create FMEs, and after computations are completed, stores results and gives access to visualization methods.
+  • "ForwardMarginalEffect": The centerpiece class of the package.
+It keeps access to a Predictor, stores important information to create FMEs, and after the computations are completed, stores results and gives access to visualization methods.
 For convenience, the wrapper function fme() can be used.
-  • "FMEPlot": An abstract class for code decoupling of different plot categories into distinct classes. Subclasses include "FMEPlotUnivariate", "FMEPlotBivariate", "FMEPlotCategorical".
-  • "ExtrapolationDetector": Identifies (and excludes) extrapolation points.
-The current implementation supports the method “envelope”, which excludes points outside of the multivariate envelope of the training data (Scholbeck et al., 2022).
-  • "NonLinearityMeasure": For the NLM, we need to approximate three line integrals, for instance via Simpson’s 3/8 rule.
-The general definition for a univariate function f(x) and integration interval $[a,b]$ corresponds to:
+  • "FMEPlot": An abstract class for code decoupling of different plot categories into distinct classes. Subclasses include "FMEPlotUnivariate", "FMEPlotBivariate", "FMEPlotHigherOrder",
+    "FMEPlotCategorical".
+  • "ExtrapolationDetector": Identifies (and excludes) EPs. The current implementation supports the method “envelope”, excluding points outside the multivariate envelope of the training data.
+  • "NonLinearityMeasure": For the NLM, we need to approximate three line integrals, e.g., via Simpson’s 3/8 rule.
+The general definition of Simpson’s 3/8 rule for a univariate function f(x) and integration interval $[a,b]$ corresponds to:
     $\int_{a}^{b}f(x)\approx\frac{b-a}{8}\left[f(a)+3f\left(\frac{2a+b}{3}\right)+3f\left(\frac{a+2b}{3}\right)+f(b)\right]$ (3)
     We make use of a composite Simpson rule, which divides up the interval $\left[a,b\right]$ into $n$ subintervals of equal size and approximates each subinterval with Eq. (3).
-  • "Partitioning": An abstract class, allowing various implementations of finding subgroups for cAMEs.
+  • "Partitioning": An abstract class, allowing for various implementations of finding subgroups for cAMEs.
 For convenience, the wrapper function came() can be used.
 The current implementation supports RP via the rpart and partykit (CTREE algorithm) packages (classes "PartitioningRPart" and "PartitioningCTREE").
-We believe there are two criteria that should guide this process:
-    • FME homogeneity inside each subgroup.
-    • The number of subgroups.
+    We believe there are two criteria that should guide this process: FME homogeneity within each subgroup and the number of subgroups.
 A low number of subgroups is generally preferred.
 In certain applications, we may want to search for a predefined number of subgroups, akin to the search for a predefined number of clusters in clustering problems.
-Many RP algorithms do not support the search for a number of subgroups, which is where the "Pruner" class comes into play.
+Many RP algorithms do not support searching for a number of subgroups, which is what the "Pruner" class is intended for.
   • "Pruner": To receive a predefined number of subgroups for arbitrary RP algorithms, we follow a two-stage process: grow a large tree by tweaking tree-specific hyperparameters and then prune it back to receive the desired number of subgroups.
-A "Partitioning" subclass is implemented such that it can first grow a large tree, for instance, with a low complexity parameter for rpart.
-Then, "Pruner" iteratively computes the SD of FMEs for each parent node of the current terminal nodes and removes all terminal nodes of the parent with the lowest SD.
-  • "PartitioningPlot": Decouples visualizations of separating $\mathcal{D}$ into subgroups from the concrete implementation of the "Partitioning" subclass.
+A "Partitioning" subclass is implemented such that it can first grow a large tree, e.g., with a low complexity parameter for rpart.
+Then "Pruner" iteratively computes the SD of FMEs for each parent node of the current terminal nodes and removes all terminal nodes of the parent with the lowest SD.
+  • "PartitioningPlot": Decouples visualizations of the separation of $\mathcal{D}$ into subgroups from specific implementations of the "Partitioning" subclass.
 Here, we make use of a dependency on partykit for a tree data structure.
@@
 Conclusion
-This paper introduces the R package fmeffects, the first software implementation of FMEs.
+This paper introduces the R package fmeffects, the first software implementation of the theory surrounding FMEs.
 We showcase the package functionality with an applied use case and discuss design choices and implications for future extensions.
-FMEs are a novel and versatile method for model-agnostic interpretations.
-They give us explanations in the form of: If we change $x$ by an amount $h$, what is the change in predicted outcome $\widehat{y}$?
-FMEs provide comprehensible and actionable advice for stakeholders, including ones without machine learning expertise.
-We therefore hope that this package will work towards a more widespread adaption and usage of FMEs in practice.
+FMEs are a versatile model-agnostic interpretation method and give us comprehensible model explanations in the form of: if we change $x$ by an amount $h$, what is the change in predicted outcome $\widehat{y}$?
+FMEs equip stakeholders, including those without ML expertise, with the ability to understand feature effects for any model.
+We therefore hope that this package will work towards a more widespread adoption of FMEs in practice.
+Software development is an ongoing process.
+As the theory surrounding FMEs evolves, so should the fmeffects package.
+As noted by Scholbeck et al.
+(2024), possible directions for future research include the development of techniques to better quantify extrapolation risk for the selection of step sizes; furthermore, the subgroup search for cAMEs is subject to uncertainties, which may be able to be quantified; and lastly, we may be able to spare computations by searching for representative FMEs, similar to prototype observations that are representative of clusters of observations (Tan et al., 2019).
+Future performance improvements may also be made via parallel computing, which at this point is only implemented for NLM computations.